# 22affiliationtext: Booth School of Business

### v1

@@ D.1.3 Regularization (MNIST)
Table 9: Latent space regularization (as in $\beta$-VAE) on MNIST
Metric / Evaluation $\beta$=0.5 $\beta$=2 $\beta$=4
Validation Metrics
Validation ELBO (bpd) ↓ 0.2591 (0.0057) 0.3399 (0.0080) 0.3725 (0.0076)
Reconstruction Loss (bpd) ↓ 0.2367 (0.0059) 0.3016 (0.0094) 0.3317 (0.0083)
Reconstructed Distribution Evaluation
Covariance Error ↓ 1.1730 (0.0739) 1.7058 (0.0478) 1.8174 (0.1025)
Mean Error ↓ 0.0180 (0.0026) 0.0232 (0.0018) 0.0149 (0.0016)
MMD ↓ 0.0031 (0.0017) 0.0268 (0.0033) 0.0688 (0.0105)
Covariance Error (ivn) ↓ 1.9956 (0.2056) 2.3965 (0.2786) 2.6715 (0.3147)
Mean Error (ivn) ↓ 0.0258 (0.0017) 0.0359 (0.0038) 0.0408 (0.0053)
MMD (ivn) ↓ 0.0821 (0.0307) 0.1110 (0.0357) 0.1491 (0.0324)
Generated Distribution Evaluation
MMD ↓ 0.0038 (0.0001) 0.0132 (0.0014) 0.0437 (0.0086)
MMD (ivn) ↓ 0.0160 (0.0041) 0.0410 (0.0112) 0.0898 (0.0179)
D.1.4 Expressivity (MNIST)
[…]
Metric / Evaluation (15, 22) (22, 22) (22, 5) (50, 22) (50, 5)
Validation Metrics
Validation ELBO (bpd) ↓ 0.3777 (0.0100) 0.3892 (0.0119) 0.2956 (0.0032) 0.3655 (0.0088) 0.3073 (0.0098)
Reconstruction Loss (bpd) ↓ 0.3532 (0.0095) 0.3641 (0.0123) 0.2653 (0.0034) 0.3413 (0.0083) 0.2781 (0.0098)
Reconstructed Distribution Evaluation (Observations)
MAE_Cov ↓ 1.6402 (0.1417) 1.7153 (0.1730) 1.4769 (0.0512) 1.6964 (0.1163) 1.5230 (0.1056)
MAE_Mean ↓ 0.0206 (0.0032) 0.0214 (0.0020) 0.0243 (0.0019) 0.0215 (0.0023) 0.0194 (0.0015)
MMD ↓ 0.0183 (0.0016) 0.0199 (0.0015) 0.0119 (0.0008) 0.0241 (0.0050) 0.0138 (0.0010)
Reconstructed Distribution Evaluation (Others)
MAE_Cov ↓ 2.8664 (0.3217) 2.9818 (0.3424) 2.3018 (0.2701) 2.7742 (0.3048) 2.3233 (0.2495)
MAE_Mean ↓ 0.0407 (0.0035) 0.0455 (0.0044) 0.0311 (0.0028) 0.0374 (0.0032) 0.0305 (0.0023)
MMD ↓ 0.1036 (0.0313) 0.1064 (0.0337) 0.0907 (0.0321) 0.1149 (0.0372) 0.0882 (0.0302)
Generated Distribution Evaluation
MMD ↓ 0.0094 (0.0006) 0.0091 (0.0015) 0.0061 (0.0008) 0.0110 (0.0021) 0.0070 (0.0017)
MMD (ivn) ↓ 0.0312 (0.0066) 0.0422 (0.0130) 0.0224 (0.0062) 0.0388 (0.0117) 0.0249 (0.0059)
[…]
Metric / Evaluation (15, 22) (15, 5) (22, 22) (22, 5) (50, 22) (50, 5)
Validation Metrics
Validation ELBO (bpd) ↓ 0.3423 (0.0089) 0.3147 (0.0105) 0.3779 (0.0178) 0.3027 (0.0063) 0.3702 (0.0208) 0.3160 (0.0124)
Reconstruction Loss (bpd) ↓ 0.3213 (0.0092) 0.2897 (0.0108) 0.3569 (0.0181) 0.2767 (0.0061) 0.3531 (0.0240) 0.2899 (0.0132)
Reconstructed Distribution Evaluation (Observations)
MAE_Cov ↓ 1.5155 (0.0766) 1.5477 (0.0886) 1.7862 (0.1572) 1.3584 (0.1025) 1.9985 (0.3544) 1.4030 (0.1516)
MAE_Mean ↓ 0.0141 (0.0014) 0.0178 (0.0019) 0.0201 (0.0030) 0.0148 (0.0031) 0.0210 (0.0054) 0.0157 (0.0029)
MMD ↓ 0.0251 (0.0021) 0.0180 (0.0026) 0.0282 (0.0039) 0.0137 (0.0018) 0.2226 (0.1939) 0.0139 (0.0030)
Reconstructed Distribution Evaluation (Others)
MAE_Cov ↓ 2.5968 (0.2735) 2.2916 (0.2427) 2.8177 (0.3093) 2.2404 (0.2412) 2.9055 (0.3518) 2.3451 (0.2558)
MAE_Mean ↓ 0.0359 (0.0035) 0.0311 (0.0024) 0.0418 (0.0040) 0.0298 (0.0023) 0.0428 (0.0068) 0.0304 (0.0028)
MMD ↓ 0.0937 (0.0279) 0.1036 (0.0359) 0.1112 (0.0337) 0.0987 (0.0350) 0.2750 (0.0712) 0.0946 (0.0333)
Generated Distribution Evaluation
MMD ↓ 0.0123 (0.0012) 0.0092 (0.0013) 0.0135 (0.0024) 0.0067 (0.0007) 0.2112 (0.1967) 0.0056 (0.0006)
MMD (ivn) ↓ 0.0334 (0.0076) 0.0272 (0.0069) 0.0374 (0.0082) 0.0242 (0.0064) 0.2254 (0.0721) 0.0245 (0.0069)
D.2 Additional NVAE Figures
D.2.1 Additional examples of concept composition

### v2

@@ E.1.3 Regularization (MNIST)
Table 8: Evaluation of CM (end-to-end) in terms of reconstruction and concept learning on MNIST for different latent space regularization strengths, as $\beta$ varies.
0.5 1.0 2.0 4.0
reconstruction (ELBO BPD)
in-distribution 0.222 ± 0.002 0.259 ± 0.005 0.298 ± 0.003 0.341 ± 0.007
out-of-distribution
concept learning (sliced Wasserstein)
obs 0.085 ± 0.003 0.099 ± 0.006 0.102 ± 0.003 0.108 ± 0.005
scaled 0.077 ± 0.006 0.062 ± 0.002 0.077 ± 0.004 0.098 ± 0.009
shear 0.108 ± 0.003 0.119 ± 0.006 0.125 ± 0.005 0.117 ± 0.004
shift 0.105 ± 0.003 0.113 ± 0.005 0.125 ± 0.004 0.130 ± 0.007
swel 0.110 ± 0.007 0.126 ± 0.007 0.129 ± 0.006 0.134 ± 0.004
thic 0.182 ± 0.011 0.178 ± 0.007 0.190 ± 0.016 0.192 ± 0.008
thin 0.069 ± 0.006 0.064 ± 0.004 0.074 ± 0.005 0.117 ± 0.005
concept learning mean 0.105 ± 0.006 0.109 ± 0.006 0.118 ± 0.006 0.128 ± 0.006
E.1.4 Regularization (quad)
This is like Appendix E.1.3 but applied to the quad dataset rather than MNIST. A description of the architecture and experiment is given in Appendix D. Tables 9 and 10 respectively show results of group lasso (Yuan and Lin, 2006) and $L_{2}$ regularization, while Table 11 shows the effects regularizing the latent space of the VAE by varying $\beta$ (Higgins et al., 2017). [^2] All other quad experiments in Appendix E.1 are run with $\lambda=0$ for group lasso and $L_{2}$ and with $\beta=1$.
Table 9: Evaluation of CM (end-to-end) in terms of reconstruction, concept learning, and composition on quad (independent) for different group norm regularization strengths.
$\lambda=$ 0 $\lambda=$1 $\lambda=$10 $\lambda=$100
reconstruction (ELBO BPD)
in-distribution 0.519 ± 0.022 0.511 ± 0.023 0.528 ± 0.030 0.492 ± 0.009
out-of-distribution 0.720 ± 0.069 0.703 ± 0.052 0.868 ± 0.147 0.993 ± 0.181
concept learning (sliced Wasserstein)
obs 0.056 ± 0.003 0.049 ± 0.001 0.054 ± 0.004 0.055 ± 0.002
quad1 0.062 ± 0.003 0.058 ± 0.004 0.056 ± 0.005 0.060 ± 0.001
quad2 0.061 ± 0.005 0.057 ± 0.006 0.057 ± 0.003 0.055 ± 0.004
quad3 0.063 ± 0.002 0.058 ± 0.004 0.058 ± 0.003 0.061 ± 0.006
quad4 0.065 ± 0.007 0.054 ± 0.003 0.052 ± 0.003 0.056 ± 0.003
size 0.075 ± 0.007 0.076 ± 0.007 0.073 ± 0.007 0.069 ± 0.005
orientation 0.057 ± 0.003 0.057 ± 0.003 0.055 ± 0.006 0.065 ± 0.006
concept learning mean 0.063 ± 0.004 0.058 ± 0.004 0.058 ± 0.004 0.060 ± 0.004
composition (sliced Wasserstein)
(quad1, quad2) 0.100 ± 0.028 0.061 ± 0.008 0.092 ± 0.023 0.101 ± 0.038
(quad1, quad3) 0.078 ± 0.016 0.080 ± 0.019 0.101 ± 0.017 0.134 ± 0.019
(quad1, quad4) 0.094 ± 0.032 0.081 ± 0.017 0.127 ± 0.040 0.132 ± 0.043
(quad1, size) 0.101 ± 0.010 0.098 ± 0.006 0.115 ± 0.011 0.127 ± 0.025
(quad1, orientation) 0.060 ± 0.003 0.055 ± 0.004 0.057 ± 0.004 0.063 ± 0.005
(quad2, quad3) 0.067 ± 0.007 0.072 ± 0.008 0.075 ± 0.008 0.123 ± 0.042
(quad2, quad4) 0.068 ± 0.004 0.087 ± 0.029 0.075 ± 0.006 0.145 ± 0.024
(quad2, size) 0.099 ± 0.008 0.097 ± 0.004 0.111 ± 0.019 0.108 ± 0.022
(quad2, orientation) 0.063 ± 0.004 0.059 ± 0.004 0.058 ± 0.004 0.065 ± 0.008
composition mean 0.081 ± 0.013 0.077 ± 0.011 0.090 ± 0.015 0.111 ± 0.025
Table 10: Evaluation of CM (end-to-end) in terms of reconstruction, concept learning, and composition on quad (independent) for different $L_{2}$-norm regularization strengths.
$\lambda=$ 0 $\lambda=$1 $\lambda=$10 $\lambda=$100
reconstruction (ELBO BPD)
in-distribution 0.519 ± 0.022 0.529 ± 0.029 0.550 ± 0.029 0.507 ± 0.024
out-of-distribution 0.720 ± 0.069 0.662 ± 0.058 0.676 ± 0.023 0.685 ± 0.040
concept learning (sliced Wasserstein)
obs 0.056 ± 0.003 0.054 ± 0.001 0.055 ± 0.002 0.054 ± 0.001
quad1 0.062 ± 0.003 0.058 ± 0.003 0.061 ± 0.003 0.053 ± 0.004
quad2 0.061 ± 0.005 0.054 ± 0.003 0.062 ± 0.002 0.056 ± 0.002
quad3 0.063 ± 0.002 0.057 ± 0.002 0.059 ± 0.005 0.055 ± 0.005
quad4 0.065 ± 0.007 0.053 ± 0.003 0.058 ± 0.002 0.052 ± 0.004
size 0.075 ± 0.007 0.073 ± 0.009 0.076 ± 0.007 0.070 ± 0.004
orientation 0.057 ± 0.003 0.058 ± 0.004 0.063 ± 0.004 0.056 ± 0.005
concept learning mean 0.063 ± 0.004 0.058 ± 0.003 0.062 ± 0.003 0.057 ± 0.004
composition (sliced Wasserstein)
(quad1, quad2) 0.100 ± 0.028 0.070 ± 0.007 0.069 ± 0.003 0.058 ± 0.003
(quad1, quad3) 0.078 ± 0.016 0.073 ± 0.014 0.096 ± 0.021 0.063 ± 0.005
(quad1, quad4) 0.094 ± 0.032 0.101 ± 0.029 0.081 ± 0.012 0.079 ± 0.019
(quad1, size) 0.101 ± 0.010 0.095 ± 0.011 0.098 ± 0.004 0.093 ± 0.007
(quad1, orientation) 0.060 ± 0.003 0.056 ± 0.004 0.061 ± 0.002 0.058 ± 0.003
(quad2, quad3) 0.067 ± 0.007 0.071 ± 0.009 0.080 ± 0.006 0.071 ± 0.009
(quad2, quad4) 0.068 ± 0.004 0.072 ± 0.009 0.067 ± 0.003 0.063 ± 0.006
(quad2, size) 0.099 ± 0.008 0.084 ± 0.006 0.097 ± 0.008 0.089 ± 0.005
(quad2, orientation) 0.063 ± 0.004 0.058 ± 0.003 0.061 ± 0.003 0.061 ± 0.003
composition mean 0.081 ± 0.013 0.076 ± 0.010 0.079 ± 0.007 0.071 ± 0.007
Table 11: Evaluation of CM (end-to-end) in terms of reconstruction, concept learning, and composition on quad (independent) for different latent space regularization strengths, as $\beta$ varies.
0.5 1.0 2.0 4.0
reconstruction (ELBO BPD)
in-distribution 0.474 ± 0.003 0.519 ± 0.022 0.599 ± 0.020 0.640 ± 0.029
out-of-distribution 0.632 ± 0.028 0.720 ± 0.069 0.821 ± 0.047 0.865 ± 0.061
concept learning (sliced Wasserstein)
obs 0.050 ± 0.001 0.056 ± 0.003 0.048 ± 0.002 0.054 ± 0.004
quad1 0.051 ± 0.003 0.062 ± 0.003 0.055 ± 0.002 0.064 ± 0.005
quad2 0.052 ± 0.002 0.061 ± 0.005 0.052 ± 0.003 0.064 ± 0.011
quad3 0.059 ± 0.002 0.063 ± 0.002 0.047 ± 0.002 0.064 ± 0.009
quad4 0.054 ± 0.002 0.065 ± 0.007 0.050 ± 0.002 0.058 ± 0.006
size 0.061 ± 0.003 0.075 ± 0.007 0.066 ± 0.005 0.082 ± 0.005
orientation 0.050 ± 0.002 0.057 ± 0.003 0.051 ± 0.002 0.068 ± 0.008
concept learning mean 0.054 ± 0.002 0.063 ± 0.004 0.053 ± 0.003 0.065 ± 0.007
composition (sliced Wasserstein)
(quad1, quad2) 0.059 ± 0.005 0.100 ± 0.028 0.080 ± 0.005 0.082 ± 0.015
(quad1, quad3) 0.070 ± 0.008 0.078 ± 0.016 0.071 ± 0.007 0.082 ± 0.014
(quad1, quad4) 0.076 ± 0.010 0.094 ± 0.032 0.073 ± 0.012 0.123 ± 0.027
(quad1, size) 0.099 ± 0.007 0.101 ± 0.010 0.118 ± 0.008 0.115 ± 0.010
(quad1, orientation) 0.052 ± 0.002 0.060 ± 0.003 0.057 ± 0.003 0.073 ± 0.005
(quad2, quad3) 0.056 ± 0.000 0.067 ± 0.007 0.074 ± 0.008 0.094 ± 0.021
(quad2, quad4) 0.087 ± 0.018 0.068 ± 0.004 0.101 ± 0.027 0.067 ± 0.009
(quad2, size) 0.086 ± 0.015 0.099 ± 0.008 0.121 ± 0.018 0.131 ± 0.011
(quad2, orientation) 0.055 ± 0.003 0.063 ± 0.004 0.056 ± 0.006 0.068 ± 0.009
composition mean 0.071 ± 0.008 0.081 ± 0.013 0.083 ± 0.010 0.093 ± 0.013
E.1.5 Expressivity (MNIST)
A description of the architecture and experiment is given in Appendix D. Tables 12 and 13 show varying widths for an expressive layer depth respectively of 2 and 5: the tuples in the column headers indicate $(0pt_{\mathrm{exp}},\ 0pt_{c})$. [^3] All other experiments on MNIST in Appendix E.1 are run with $h_{\mathrm{exp}}=2$, $0pt_{\mathrm{exp}}=15$, and $0pt_{c}=5$. These results suggest relatively stable results across different configurations for concept learning while the best reconstruction performance is obtained from a moderate bottleneck, with expressive layer input width $0pt_{\mathrm{exp}}=22$ and a concept width of $0pt_{c}=5$.
Table 12: Evaluation of CM (end-to-end) in terms of reconstruction and concept learning on MNIST for different capacities, with depth $h_{exp}=2$ as $(w_{exp},w_{c})$ varies.
$(w_{exp},w_{c})$ (15, 22) (15, 5) (22, 22) (22, 5) (50, 22) (50, 5)
reconstruction (ELBO BPD)
in-distribution 0.260 ± 0.005 0.259 ± 0.005 0.255 ± 0.003 0.252 ± 0.003 0.265 ± 0.001 0.265 ± 0.006
out-of-distribution
concept learning (sliced Wasserstein)
obs 0.087 ± 0.003 0.099 ± 0.006 0.082 ± 0.005 0.095 ± 0.006 0.086 ± 0.004 0.092 ± 0.008
scaled 0.073 ± 0.011 0.062 ± 0.002 0.065 ± 0.004 0.067 ± 0.005 0.068 ± 0.003 0.072 ± 0.004
shear 0.115 ± 0.006 0.119 ± 0.006 0.108 ± 0.006 0.122 ± 0.006 0.116 ± 0.004 0.118 ± 0.004
shift 0.106 ± 0.009 0.113 ± 0.005 0.108 ± 0.008 0.111 ± 0.002 0.096 ± 0.003 0.115 ± 0.005
swel 0.119 ± 0.006 0.126 ± 0.007 0.107 ± 0.005 0.118 ± 0.006 0.111 ± 0.007 0.118 ± 0.007
thic 0.202 ± 0.016 0.178 ± 0.007 0.180 ± 0.007 0.171 ± 0.014 0.176 ± 0.010 0.178 ± 0.015
thin 0.065 ± 0.006 0.064 ± 0.004 0.069 ± 0.005 0.072 ± 0.010 0.073 ± 0.007 0.067 ± 0.006
concept learning mean 0.110 ± 0.008 0.109 ± 0.006 0.103 ± 0.006 0.108 ± 0.007 0.104 ± 0.005 0.109 ± 0.007
Table 13: Evaluation of CM (end-to-end) in terms of reconstruction and concept learning on MNIST for different capacities, with depth $h_{exp}=4$ as $(w_{exp},w_{c})$ varies.
$(w_{exp},w_{c})$ (15, 22) (15, 5) (22, 22) (22, 5) (50, 22) (50, 5)
reconstruction (ELBO BPD)
in-distribution 0.269 ± 0.003 0.272 ± 0.006 0.276 ± 0.004 0.255 ± 0.004 0.281 ± 0.005 0.279 ± 0.008
out-of-distribution
concept learning (sliced Wasserstein)
obs 0.096 ± 0.004 0.088 ± 0.004 0.089 ± 0.001 0.092 ± 0.004 0.090 ± 0.005 0.093 ± 0.003
scaled 0.063 ± 0.003 0.071 ± 0.004 0.071 ± 0.003 0.085 ± 0.018 0.067 ± 0.001 0.079 ± 0.003
shear 0.119 ± 0.006 0.110 ± 0.004 0.116 ± 0.004 0.113 ± 0.004 0.107 ± 0.004 0.103 ± 0.004
shift 0.114 ± 0.004 0.109 ± 0.006 0.108 ± 0.005 0.107 ± 0.005 0.113 ± 0.006 0.104 ± 0.002
swel 0.120 ± 0.006 0.113 ± 0.007 0.110 ± 0.001 0.122 ± 0.006 0.116 ± 0.005 0.115 ± 0.003
thic 0.179 ± 0.012 0.178 ± 0.014 0.175 ± 0.011 0.185 ± 0.011 0.188 ± 0.007 0.184 ± 0.011
thin 0.066 ± 0.005 0.076 ± 0.004 0.072 ± 0.003 0.070 ± 0.003 0.076 ± 0.007 0.083 ± 0.006
concept learning mean 0.108 ± 0.005 0.106 ± 0.006 0.106 ± 0.004 0.111 ± 0.007 0.108 ± 0.005 0.109 ± 0.005
E.1.6 Expressivity (quad)
This is like Appendix E.1.5 but applied to the quad dataset rather than MNIST. A description of the architecture and experiment is given in Appendix D. Tables 14 and 15 show varying widths for an expressive layer depth respectively of 2 and 5: the tuples in the column headers indicate $(0pt_{\mathrm{exp}},\ 0pt_{c})$. [^4] All other experiments on quad in Appendix E.1 are run with $h_{\mathrm{exp}}=2$, $0pt_{\mathrm{exp}}=32$, and $0pt_{c}=8$. These results suggest relatively stable results across different configurations.
Table 14: Evaluation of CM (end-to-end) in terms of reconstruction, concept learning, and composition on quad (independent) for different capacities, with depth $h_{exp}=2$ as $(w_{exp},w_{c})$ varies.
$(w_{exp)},w_{c)})$ (16, 16) (16, 32) (16, 8) (32, 16) (32, 32) (32, 8)
reconstruction (ELBO BPD)
in-distribution 0.494 ± 0.014 0.497 ± 0.012 0.494 ± 0.010 0.512 ± 0.023 0.557 ± 0.016 0.519 ± 0.022
out-of-distribution 0.676 ± 0.024 0.665 ± 0.025 0.598 ± 0.025 0.626 ± 0.030 0.752 ± 0.067 0.720 ± 0.069
concept learning (sliced Wasserstein)
obs 0.055 ± 0.005 0.056 ± 0.002 0.056 ± 0.004 0.053 ± 0.003 0.063 ± 0.005 0.056 ± 0.003
quad1 0.057 ± 0.004 0.061 ± 0.005 0.059 ± 0.004 0.058 ± 0.004 0.067 ± 0.007 0.062 ± 0.003
quad2 0.061 ± 0.006 0.053 ± 0.003 0.060 ± 0.003 0.056 ± 0.005 0.063 ± 0.005 0.061 ± 0.005
quad3 0.058 ± 0.004 0.058 ± 0.005 0.065 ± 0.007 0.055 ± 0.003 0.066 ± 0.006 0.063 ± 0.002
quad4 0.060 ± 0.007 0.063 ± 0.004 0.063 ± 0.006 0.055 ± 0.003 0.068 ± 0.006 0.065 ± 0.007
size 0.072 ± 0.008 0.072 ± 0.007 0.076 ± 0.005 0.071 ± 0.005 0.082 ± 0.011 0.075 ± 0.007
orientation 0.061 ± 0.006 0.055 ± 0.002 0.058 ± 0.003 0.055 ± 0.003 0.062 ± 0.006 0.057 ± 0.003
concept learning mean 0.061 ± 0.006 0.060 ± 0.004 0.062 ± 0.005 0.057 ± 0.004 0.067 ± 0.007 0.063 ± 0.004
composition (sliced Wasserstein)
(quad1, quad2) 0.091 ± 0.010 0.086 ± 0.010 0.066 ± 0.006 0.076 ± 0.014 0.131 ± 0.017 0.100 ± 0.028
(quad1, quad3) 0.064 ± 0.008 0.083 ± 0.006 0.075 ± 0.012 0.059 ± 0.005 0.102 ± 0.010 0.078 ± 0.016
(quad1, quad4) 0.082 ± 0.019 0.069 ± 0.006 0.064 ± 0.007 0.076 ± 0.012 0.092 ± 0.013 0.094 ± 0.032
(quad1, size) 0.084 ± 0.007 0.098 ± 0.009 0.094 ± 0.005 0.107 ± 0.013 0.117 ± 0.011 0.101 ± 0.010
(quad1, orientation) 0.059 ± 0.004 0.065 ± 0.006 0.060 ± 0.004 0.055 ± 0.003 0.068 ± 0.011 0.060 ± 0.003
(quad2, quad3) 0.083 ± 0.017 0.080 ± 0.014 0.076 ± 0.015 0.066 ± 0.008 0.083 ± 0.012 0.067 ± 0.007
(quad2, quad4) 0.079 ± 0.010 0.074 ± 0.007 0.071 ± 0.010 0.065 ± 0.009 0.077 ± 0.008 0.068 ± 0.004
(quad2, size) 0.100 ± 0.014 0.085 ± 0.011 0.087 ± 0.006 0.093 ± 0.006 0.116 ± 0.011 0.099 ± 0.008
(quad2, orientation) 0.062 ± 0.006 0.057 ± 0.003 0.060 ± 0.005 0.057 ± 0.003 0.067 ± 0.008 0.063 ± 0.004
composition mean 0.078 ± 0.011 0.077 ± 0.008 0.072 ± 0.008 0.073 ± 0.008 0.095 ± 0.011 0.081 ± 0.013
Table 15: Evaluation of CM (end-to-end) in terms of reconstruction, concept learning, and composition on quad (independent) for different capacities, with depth $h_{exp}=4$ as $(w_{exp},w_{c})$ varies.
$(w_{exp},w_{c})$ (16, 16) (16, 32) (16, 8) (32, 16) (32, 32) (32, 8)
reconstruction (ELBO BPD)
in-distribution 0.511 ± 0.008 0.516 ± 0.015 0.573 ± 0.045 0.590 ± 0.009 0.613 ± 0.020 0.580 ± 0.017
out-of-distribution 0.761 ± 0.040 0.750 ± 0.033 0.674 ± 0.049 0.706 ± 0.035 0.759 ± 0.081 0.636 ± 0.017
concept learning (sliced Wasserstein)
obs 0.059 ± 0.004 0.052 ± 0.002 0.052 ± 0.003 0.053 ± 0.002 0.059 ± 0.006 0.065 ± 0.004
quad1 0.061 ± 0.005 0.054 ± 0.004 0.061 ± 0.005 0.063 ± 0.002 0.081 ± 0.013 0.075 ± 0.005
quad2 0.057 ± 0.004 0.057 ± 0.002 0.058 ± 0.003 0.064 ± 0.004 0.080 ± 0.014 0.071 ± 0.005
quad3 0.059 ± 0.006 0.058 ± 0.003 0.064 ± 0.003 0.066 ± 0.004 0.078 ± 0.014 0.073 ± 0.002
quad4 0.056 ± 0.004 0.053 ± 0.002 0.064 ± 0.003 0.058 ± 0.004 0.075 ± 0.012 0.071 ± 0.002
size 0.069 ± 0.003 0.081 ± 0.005 0.077 ± 0.008 0.097 ± 0.004 0.111 ± 0.016 0.087 ± 0.006
orientation 0.059 ± 0.004 0.055 ± 0.003 0.061 ± 0.005 0.060 ± 0.003 0.071 ± 0.012 0.071 ± 0.003
concept learning mean 0.060 ± 0.004 0.059 ± 0.003 0.063 ± 0.004 0.066 ± 0.003 0.079 ± 0.012 0.073 ± 0.004
composition (sliced Wasserstein)
(quad1, quad2) 0.071 ± 0.008 0.090 ± 0.010 0.059 ± 0.004 0.089 ± 0.009 0.091 ± 0.016 0.074 ± 0.005
(quad1, quad3) 0.062 ± 0.006 0.062 ± 0.004 0.080 ± 0.023 0.075 ± 0.005 0.086 ± 0.014 0.079 ± 0.006
(quad1, quad4) 0.112 ± 0.025 0.072 ± 0.010 0.063 ± 0.009 0.077 ± 0.004 0.086 ± 0.010 0.083 ± 0.005
(quad1, size) 0.083 ± 0.003 0.119 ± 0.008 0.102 ± 0.007 0.106 ± 0.003 0.133 ± 0.008 0.112 ± 0.005
(quad1, orientation) 0.059 ± 0.003 0.057 ± 0.003 0.062 ± 0.006 0.064 ± 0.003 0.079 ± 0.011 0.074 ± 0.004
(quad2, quad3) 0.077 ± 0.011 0.072 ± 0.002 0.061 ± 0.004 0.083 ± 0.008 0.109 ± 0.028 0.076 ± 0.006
(quad2, quad4) 0.092 ± 0.016 0.083 ± 0.017 0.065 ± 0.004 0.081 ± 0.005 0.102 ± 0.021 0.075 ± 0.007
(quad2, size) 0.091 ± 0.006 0.108 ± 0.009 0.101 ± 0.016 0.128 ± 0.012 0.135 ± 0.021 0.108 ± 0.007
(quad2, orientation) 0.058 ± 0.004 0.062 ± 0.003 0.064 ± 0.002 0.067 ± 0.005 0.081 ± 0.015 0.073 ± 0.005
composition mean 0.078 ± 0.009 0.080 ± 0.007 0.073 ± 0.008 0.086 ± 0.006 0.100 ± 0.016 0.084 ± 0.005
E.2 Additional NVAE Results
E.2.1 Additional examples of concept composition